# Toward Routing River Water in Land Surface Models with Recurrent Neural Networks

### v1

@@ Acknowledgements.
This research was supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program. We thank Frederik Kratzert for a helpful early conversation, as well as the other maintainers of neuralhydrology, for the clear and helpful code. We also acknowledge Akshay Sridhar, who assisted with technical parts of the code and data handling, as well as insightful conversations. All numerical calculations and model training used for this manuscript were performed with the help of Caltech’s Resnick High Performance Computing Center.References
@@
- GRDC (2023) GRDC: Global Runoff Data Center, https://www.bafg.de/GRDC/EN/Home/homepage_node.html, accessed on July 7, 2023, 2023.
@@
- Harrigan et al. (2020) Harrigan, S., Zsoter, E., Alfieri, L., Prudhomme, C., Salamon, P., Wetterhall, F., Barnard, C., Cloke, H., and Pappenberger, F.: GloFAS-ERA5 operational global river discharge reanalysis 1979–present, Earth System Science Data, 12, 2043–2060, 10.5194/essd-12-2043-2020, 2020.
@@
- Lima (2024) Lima, M.: Rivers, 10.5281/zenodo.10975804, 2024.

### v3

@@ Acknowledgements.
This research was supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program. We thank Frederik Kratzert for a helpful early conversation, as well as the other maintainers of neuralhydrology, for the clear and helpful code. We thank Svea Bertolatus, Stefania Grimaldi and the GloFAS team for providing the set of gauges from GRDC used to calibrate the model that was used to download the discharge reanalysis data, as well as helpful clarifications about GloFAS versions. ML acknowledges Akshay Sridhar, who assisted with technical parts of the code, data handling and insightful conversations, as well as Thomas Dubos for helping this research project to take place. All numerical calculations and model training used for this manuscript were performed with the help of Caltech’s Resnick High Performance Computing Center.References
- Addor et al. (2017) Addor, N., Newman, A. J., Mizukami, N., and Clark, M. P.: The CAMELS data set: catchment attributes and meteorology for large-sample studies, Hydrology and Earth System Sciences, 21, 5293–5313, 10.5194/hess-21-5293-2017, 2017.
@@
- Balsamo et al. (2009) Balsamo, G., Beljaars, A., Scipal, K., Viterbo, P., Van Den Hurk, B., Hirschi, M., and Betts, A. K.: A Revised Hydrology for the ECMWF Model: Verification from Field Site to Terrestrial Water Storage and Impact in the Integrated Forecast System, Journal of Hydrometeorology, 10, 623–643, 10.1175/2008jhm1068.1, 2009.
@@
- Branstetter (2001) Branstetter, M. L.: Development of a parallel river transport algorithm and applications to climate studies, The University of Texas at Austin, 2001.
@@
- Hersbach et al. (2020) Hersbach, H., Bell, B., Berrisford, P., Hirahara, S., Horányi, A., Muñoz-Sabater, J., Nicolas, J., Peubey, C., Radu, R., Schepers, D., Simmons, A., Soci, C., Abdalla, S., Abellan, X., Balsamo, G., Bechtold, P., Biavati, G., Bidlot, J., Bonavita, M., De Chiara, G., Dahlgren, P., Dee, D., Diamantakis, M., Dragani, R., Flemming, J., Forbes, R., Fuentes, M., Geer, A., Haimberger, L., Healy, S., Hogan, R. J., Hólm, E., Janisková, M., Keeley, S., Laloyaux, P., Lopez, P., Lupu, C., Radnoti, G., De Rosnay, P., Rozum, I., Vamborg, F., Villaume, S., and Thépaut, J.: The ERA5 global reanalysis, Quarterly Journal of the Royal Meteorological Society, 146, 1999–2049, 10.1002/qj.3803, 2020.
@@
- Kingma and Ba (2014) Kingma, D. P. and Ba, J.: Adam: A method for stochastic optimization, URL https://arxiv.org/abs/1412.6980, 2014.
@@
- Lima (2024) Lima, M.: Rivers, 10.5281/zenodo.13752864, 2024.
@@
- Lohmann et al. (1996) Lohmann, D., Nolte-Holube, R., and Raschke, E.: A large-scale horizontal routing model to be coupled to land surface parametrization schemes, Tellus A Dynamic Meteorology and Oceanography, 48, 708–721, 10.1034/j.1600-0870.1996.t01-3-00009.x, 1996.
@@
- Moshe et al. (2020) Moshe, Z., Metzger, A., Elidan, G., Kratzert, F., Nevo, S., and El-Yaniv, R.: HydroNeTs: Leveraging river structure for hydrologic modeling, URL https://arxiv.org/abs/2007.00595, 2020.
@@
- Oki and Sud (1998) Oki, T. and Sud, Y. C.: Design of Total Runoff Integrating Pathways (TRIP)—A global river channel network, AMETSOC, 10.1175/1087-3562(1998)002, 1998.
@@
- Wu et al. (2014) Wu, H., Adler, R. F., Tian, Y., Huffman, G. J., Li, H., and Wang, J.: Real-time global flood estimation using satellite-based precipitation and a coupled land surface and routing model, Water Resources Research, 50, 2693–2717, 10.1002/2013wr014710, 2014.
- Ye et al. (2012) Ye, A., Duan, Q., Zhan, C., Liu, Z., and Mao, Y.: Improving kinematic wave routing scheme in Community Land Model, Hydrology Research, 44, 886–903, 10.2166/nh.2012.145, 2012.
- Zomer et al. (2008) Zomer, R. J., Trabucco, A., Bossio, D. A., and Verchot, L.: Climate change mitigation: A spatial analysis of global land suitability for clean development mechanism afforestation and reforestation, Agriculture, Ecosystems & Environment, 126, 67–80, 10.1016/j.agee.2008.01.014, 2008.
- Zomer et al. (2022) Zomer, R. J., Xu, J., and Trabucco, A.: Version 3 of the Global Aridity Index and Potential Evapotranspiration Database, Scientific Data, 9, 10.1038/s41597-022-01493-1, 2022.